# Cosmic dynamics and observational constraints in $f(Q)$ gravity with affine equation of state

### v1

@@ 5.3 Age of the Universe
The age of universe is an important criterion to test the viability of model as compared to the observable universe. The cosmic age $t(z)$ may be expressed in terms of the red-shift $z$ as  [92]
@@
In model II, the universe present age value is $t_{0}=13.58^{+0.33}_{-0.22}$ Gyr and $t_{0}=13.59^{+0.69}_{-0.32}$ Gyr for CC and joint estimates respectively, both of which are again close with the current age values of the $\Lambda$CDM model mentioned above.
@@ 6 Conclusions
We probe the possibility to describe the unified scenario of universe expansion in $f(Q)$ gravity. The late-time accelerated expansion of the Universe has been investigated in $f(Q)$ gravity framework with affine equation of state (EoS) $(p_{d}=\alpha\rho_{d}-\rho_{0})$. In the $f(Q)$ model with affine EoS, we show that the unified scenario may be realized where the non-metricity may yield the transition from the decelerated expansion into accelerated expansion. In this model, a particular functional form $f(Q)=c\sqrt{Q}$ of $f(Q)$ gravity has been derived and it is shown that the matter dominated expansion may be also be explained with the non-metricity scalar in the model. On the other hand, we also study the model composed of cold dark matter and the dark fluid satisfying the affine EoS. A descriptive differences between these models have been studied.
@@
We use the observational data of the CC and supernovae type Ia Pantheon data for the Markov Chain Monte Carlo (MCMC) analysis based on the Hubble parameters (15) and (16) of the derived models. A summary of the constrained values extracted from the MCMC application in the models are provided in Tables (1) and (2).
@@
The evolution dynamics in the present study has been examined by study of the cosmological parameters. The behavior of deceleration parameter has been investigated. The evolution and dominant cosmic components during various epochs are reflected in the deceleration parameter behaviors of both models. The universe’s phase recently transitioned from decelerated to accelerated as indicated by the deceleration parameter’s evolution curve. In Model I, the estimated present deceleration parameter values are $q_{0}=-0.56$ and $q_{0}=-0.58$ for the CC and joint analyses respectively. For Model II, we obtain the deceleration parameter values at present as $q_{0}=-0.536$ and $q_{0}=-0.548$ based on the CC and joint data respectively. A negative deceleration parameter at $z=0$ demonstrates that cosmic expansion is accelerating for both models. At late-times, the negative pressure introduced by the non-metricity may be responsible for the accelerated universe expansion.
@@
In the considered models, the physical behavior has been elaborated on the basis of energy density, pressure and EoS parameter for the model parameters’s constrained values . The energy density remains positive whereas pressure exhibits a change from positive to negative values. It is observed in the models that the cosmic accelerated expansion is caused by this negative pressure available in late era of the universe expansion. The behavior of energy density and pressure rules out the phantom evolution in the models. The varying dark energy evolution may correspond to the cosmological constant-like behavior in the asymptotic era having $z\rightarrow-1$.
@@
In particular, for the model I, at $z=0$, $\omega_{d}=-0.706$ and $\omega_{d}=-0.720$ for CC and joint data respectively. These values at the present time suggest the existence of a quintessence-type dark energy in the model. Similarly, for the Model II, the EoS parameter values are obtained $\omega_{e}=-0.690$ (for CC dataset) and $\omega_{e}=-0.698$ (for joint dataset) at $z=0$. The graphical behavior reveals that the EoS parameter for dark energy exhibits quintessence kind of dark energy. In summary, the $f(Q)$ model with affine EoS is consistent with the observations of cosmic acceleration. The null energy condition’s validity in the models prohibits the model to possess finite-time future singularities.
@@
For the $\Lambda$CDM model, the cosmological parameters are $q=-1$, $\omega=-1$, $j=1$, Snap $(s)=1$ during the asymptotic era as $z\rightarrow-1$. In the de-Sitter model of cosmic expansion, where $H$ remains constant and $q=-1$, $j=1$ and $s=1$. This indicates that the universe is undergoing exponential acceleration with the rate of change of both acceleration and jerk are increasing. In this context, for the Model I (Eq. 15), the dark energy equation of state (EoS) parameter ($\omega_{d})$ at present is near $-0.7$ and the deceleration parameter is approximately $-0.56$(for CC data) and $-0.58$(for joint data). In this scenario, the universe’s acceleration is attributed to dynamical dark energy and the jerk parameter is close to $\Lambda$CDM limit $(j=1)$ while the snap parameter is negative for the median values of model parameters. The Tables (1) and (2) highlight a summary of the results on cosmographic parameters. It is worthwhile to mentioned that these quantities will eventually align with the $\Lambda$CDM limits in the distant future. However, at present times, the cosmic dynamics of these model are differing from the $\Lambda$CDM model.
@@
The age of universe is also estimated in the context of present $f(Q)$ gravity models. For model I, the value is $t_{0}=13.51^{+0.08}_{-0.37}$ Gyr and $t_{0}=13.41^{+0.65}_{-1.04}$ Gyr for CC and joint estimates respectively. For model II, the universe’s present age value is $t_{0}=13.58^{+0.33}_{-0.22}$ Gyr (for CC data) and $t_{0}=13.59^{+0.69}_{-0.32}$ Gyr (for joint data).
@@
In summary, we show that the observationally compatible unified scenario may be realized in the $f(Q)$ gravity framework with the affine EoS. The non-metricity scalar may explain the matter dominated era of the past along with the dark energy dominated era of the observable universe. A descriptive difference between two class of models have illustrated that the present day matter density may have the lower value (as compared to $\Lambda$CDM model) in a model having cold dark matter and dark fluid satisfying affine EoS.
Acknowledgements

### v2

@@ 5.3 Age of the Universe
The age of the universe is an important criterion to test the viability of the model as compared to the observable universe. The cosmic age $t(z)$ may be expressed in terms of the red-shift $z$ as  [96]
@@
In model II, the universe present age value is $t_{0}=13.76$ Gyr and $t_{0}=13.59$ Gyr for CC and joint estimates respectively, both of which are again close with the current age values of the $\Lambda$CDM model mentioned above.
@@ 6 Conclusions
We probe the possibility to describe the unified scenario of universe expansion in $f(Q)$ gravity. The late-time accelerated expansion of the Universe has been investigated in the $f(Q)$ gravity framework with an affine equation of state (EoS) $(p_{d}=\alpha\rho_{d}-\rho_{0})$. In the $f(Q)$ model with affine EoS, we show that the unified scenario may be realized where the non-metricity may yield the transition from the decelerated expansion into accelerated expansion. In this model, a particular functional form $f(Q)=C\sqrt{Q}$ of $f(Q)$ gravity has been derived and it is shown that the matter dominated expansion may also be explained with the non-metricity scalar in the model. In other words, we show that the unified cosmological scenario yielded by the affine EoS fluid may belong to class of $f(Q)=C\sqrt{Q}$ cosmologies in the $f(Q)$ gravity framework. Further, we also study the model composed of cold dark matter and the dark fluid satisfying the affine EoS. Descriptive differences between these models have been studied.
@@
We use the observational data of the CC and supernovae type Ia Pantheon data for the MCMC analysis based on the Hubble parameters (15) and (16) of the derived models. A summary of the constrained values extracted from the MCMC application in the models are provided in Tables (1) and (2).
@@
The evolution dynamics in the present study has been examined by study of the cosmological parameters. The behavior of the deceleration parameter has been investigated. The evolution and dominant cosmic components during various epochs are reflected in the deceleration parameter behaviors of both models. The universe’s phase recently transitioned from decelerated to accelerated as indicated by the deceleration parameter’s evolution curve. In Model I, the estimated present deceleration parameter values are $q_{0}=-0.56$ and $q_{0}=-0.58$ for the CC and joint analyses respectively. For Model II, we obtain the deceleration parameter values at present as $q_{0}=-0.881$ and $q_{0}=-0.597$ based on the CC and joint data respectively. A negative deceleration parameter at $z=0$ demonstrates that cosmic expansion is accelerating for both models. At late-times, the negative pressure introduced by the non-metricity may be responsible for the accelerated universe expansion.
@@
In the considered models, the physical behavior has been elaborated on the basis of energy density, pressure and EoS parameters for the model parameters’ constrained values. The energy density remains positive whereas pressure exhibits a change from positive to negative values. It is observed in the models that the cosmic accelerated expansion is caused by this negative pressure available in the late era of the universe expansion. The universe in Model I may approach to the $\Lambda$CDM limits in future as $z\rightarrow-1$. The varying dark energy evolution may correspond to the cosmological constant-like behavior in the asymptotic era having $z\rightarrow-1$. While the energy density of the universe in Model II will increase in future due to phantom-like characteristics. The parameter constraints subjected to the CC and Pantheon data suggest that Model I is classically stable while Model II is classically unstable. It is an important distinction as far as physical acceptability of these models are concerned.
@@
In particular, for the model I, at $z=0$, $\omega_{e}=-0.706$ and $\omega_{e}=-0.720$ for CC and joint data respectively. These values at the present time suggest the existence of a quintessence-type dark energy in the model. Similarly, for the Model II, the EoS parameter values are obtained $\omega_{e}=-0.921$ (for CC dataset) and $\omega_{e}=-0.731$ (for joint dataset) at $z=0$. The graphical behavior reveals that the EoS parameter for dark energy exhibits quintessence kind of dark energy at present times. In summary, the $f(Q)$ model with affine EoS is consistent with the observations of cosmic acceleration. The violation of null energy condition in the Model II may not prohibit from the existence of finite-time future singularity.
@@
For the $\Lambda$CDM model, the cosmological parameters are $q=-1$, $\omega=-1$, $j=1$, Snap $(s)=1$ during the asymptotic era as $z\rightarrow-1$. In other words, in the de-Sitter model of cosmic expansion, where $H$ remains constant and $q=-1$, $j=1$ and $s=1$. This indicates that the universe is undergoing exponential acceleration with the rate of change of both acceleration and jerk increasing. In this context, for the classically stable Model I (Eq. 15), the dark energy equation of state (EoS) parameter ($\omega_{e})$ at present is near $-0.7$ and the deceleration parameter is approximately $-0.56$ (for CC data) and $-0.58$ (for joint data). In this scenario, the universe’s acceleration is attributed to dynamical dark energy and the jerk parameter is close to $\Lambda$CDM limit $(j=1)$ while the snap parameter is negative for the median values of model parameters. The Tables (1) and (2) highlight a summary of the results on cosmographic parameters. It is worthwhile to mention that these quantities will eventually align with the $\Lambda$CDM limits in the distant future. However, at present times, the cosmic dynamics of these models are differing from the $\Lambda$CDM model.
@@
The age of the universe is also estimated in the context of present $f(Q)$ gravity models. For model I, the value is $t_{0}=13.51$ Gyr and $t_{0}=13.41$ Gyr for CC and joint estimates respectively. For model II, the universe’s present age value is $t_{0}=13.76$ Gyr (for CC data) and $t_{0}=13.59$ Gyr (for joint data).
@@
In summary, we show that the observationally compatible unified scenario may be realized in the $f(Q)$ gravity framework with the affine EoS. The non-metricity scalar may explain the matter dominated era of the past along with the dark energy dominated era of the observable universe. A descriptive difference between two classes of models have illustrated that the present day matter density of the considered models are approximately the same but the Model I is classically stable while Model II is classically unstable. The dynamical dark energy in Model I is of quintessence kind with $\Lambda$CDM limit in future while the dynamical dark energy of Model II may have the phantom nature in future.
Acknowledgments